# Real-Time Sampling-Based Safe Motion Planning for Robotic Manipulators in Dynamic Environments

### v1

@@ Abstract
In this paper, we present the main features of Dynamic Rapidly-exploring Generalized Bur Tree (DRGBT) algorithm, a sampling-based planner for dynamic environments. We provide a detailed time analysis and appropriate scheduling to facilitate a real-time operation. To this end, an extensive analysis is conducted to identify the time-critical routines and their dependence on the number of obstacles. Furthermore, information about the distance to obstacles is used to compute a structure called dynamic expanded bubble of free configuration space, which is then utilized to establish sufficient conditions for a guaranteed safe motion of the robot while satisfying all kinematic constraints. An extensive randomized simulation trial is conducted to compare the proposed algorithm to a competing state-of-the-art method. Finally, an experimental study on a real robot is carried out covering a variety of scenarios including those with human presence. The results show the effectiveness and feasibility of real-time execution of the proposed motion planning algorithm within a typical sensor-based arrangement, using cheap hardware and sequential architecture, without the necessity for GPUs or heavy parallelization.
@@ I Introduction
TABLE I: Summary of the relevant features of the state-of-the-art algorithms
@@
alg. frequ-
ency $\mathrm{[Hz]}$
@@ 2.60 GHz CPU
RRT-based algorithms for DEs have been concurrently developed with those based on PRM. One of the first approaches is execution-extended RRT (ERRT) algorithm [14], which improves replanning efficiency and the quality of generated paths by using a so-called waypoint cache memory in $\mathcal{C}$-space. Afterwards, dynamic RRT (DRRT) algorithm [15], based on the D* family of deterministic algorithms [16], is proposed. It simply removes occupied edges in a tree, while collision-free ones are retained. Multipartite RRT (MP-RRT) algorithm [17] combines the advantages of existing RRT adaptations for dynamic MP from [18, 9, 15]. This is conveniently achieved by affecting a sampling distribution and reusing previously planned edges. Chance-constrained RRT [19] is an RTMP algorithm known for using so-called chance constraints to guarantee probabilistic feasibility in DE.
@@
Hardware acceleration has long been used in MP. The authors in [43] developed a dedicated FPGA-based circuitry to solve the planning problem in less than $\mathrm{1\,[ms]}$ for a 6-DoF manipulator. The collision detection time for PRM edges is demonstrated to be virtually independent of the graph size.
On the other hand, a massive GPU acceleration has been applied in [44] demonstrating a parallel optimization technique to solve the MP problem for manipulators. A significant speedup over state-of-the-art methods is achieved. GPU-accelerated library – CuRobo is released, which offers a parallel geometric planner and a collision-free IK solver.
@@
It is worth stressing that all of the aforementioned planners deal with one or more issues listed below:
  • planning is supported only for mobile robots or robotic manipulators with fewer than six DoFs;
@@
To this end, we propose a novel approach that aims at overcoming all of the aforementioned issues. Relevant features of the selected set of state-of-the-art algorithms are summarized within Tab. I. The sign “n/a” stands for the feature that was either not available or not possible to find despite our best efforts.
@@
  • a comprehensive randomized simulation trial revealing DRGBT outperforms a competing state-of-the-art method;
@@
The remainder of the paper is structured as follows. Sec. II provides an overview of DRGBT algorithm from [50]. Sec. III proposes a scheduling framework for the real-time execution of the algorithm. Sec. IV proposes a concept of dynamic bubbles and burs, and proves that a guaranteed safe motion of the robot in DEs can be achieved. A simulation study within Sec. V reveals the most critical routines/procedures, and confirms that the chosen scheduling framework ensures the real-time execution of the algorithm. Moreover, DRGBT is compared to a state-of-the-art method within an extensive randomized trial. Sec. VI provides experimental results demonstrating the real-time execution. Finally, Sec. VII brings some discussion, concluding remarks, and future work directions.
Figure 1: Generalized bur with 30 “spines” (radial directions) and 5 “layers” (extensions along a single spine) (right) corresponding to a configuration of 2-DoF manipulator in the workspace with four world obstacles (left) [51]
Figure 2: Graphical interpretation of some DRGBT components. (a) The obstacle (black) blocks the predefined path (gray) to the goal; (b) The obstacle has moved in a way that clears the path to the goal, thus replanning has been performed and predefined path has been changed; (c) The horizon $\boldsymbol{Q}_{h}$ and the reached horizon $\color[rgb]{1,0,0}\boldsymbol{Q}_{h}^{*}$ are shown, where the horizon spines (i.e., generalized bur) are depicted by solid red lines.
@@ II DRGBT Algorithm
For better comprehension, we first introduce the notation and define the problem statement in Subsec. II-A. Then, we broadly describe DRGBT algorithm with its main features through Subsecs. II-B and II-C, while the details including pseudocode are given in Subsec. II-D.
@@ II-A Problem Statement and Assumptions
The robot’s $n$-dimensional configuration space is denoted as $\mathcal{C}$. The obstacle space $\mathcal{C}_{obs}\subset\mathcal{C}$ is the open subset of $\mathcal{C}$ which implies the collision with obstacles or self-collision. The free space $\mathcal{C}_{free}=\mathcal{C}\backslash\mathcal{C}_{obs}$ is the closed subset of $\mathcal{C}$ that the robot can reach. It is assumed that $\mathcal{C}_{obs}$ corresponds to a finite number $N_{obs}$ of (possibly overlapping) convex world obstacles $\mathcal{WO}_{j}$, $j\in\{1,2,\dots,N_{obs}\}$, existing in the robot’s workspace $\mathcal{W}$. The case of non-convex obstacles can be tackled whether by using some explicit decomposition techniques (e.g., [52]), or indirectly through algorithms for collision/distance checking that use an implicit representation of the environment via bounding volume hierarchies (BVHs), which typically rely on convex geometric primitives [53].
@@
A motion trajectory $\boldsymbol{\pi}(t)=\boldsymbol{\pi}[\boldsymbol{q}_{0},\boldsymbol{q}_{f}]$ represents a curve defined by a continuous mapping $\boldsymbol{\pi}:t\in[t_{0},t_{f}]\mapsto\mathcal{C}$ such that $\boldsymbol{\pi}(t_{0})\mapsto\boldsymbol{q}_{0}$ and $\boldsymbol{\pi}(t_{f})\mapsto\boldsymbol{q}_{f}$, where $t_{0}$ is the initial time and $t_{f}$ is a final time (in general not a priori given). A trajectory is considered valid iff both $\boldsymbol{\pi}(t)\cap\mathcal{C}_{obs}(t)=\emptyset$, $\forall t\in[t_{0},t_{f}]$, and it is feasible for the robot to follow $\boldsymbol{\pi}(t)$ within the limits of its kinematics and other constraints denoted with $\mathcal{K}$.
@@
II-B A Brief Overview of the Algorithm
@@
DRGBT algorithm, originally proposed in [50], uses the idea of (generalized) burs of free $\mathcal{C}$-space from [54] and [51]. The bur is proven to capture relatively large portions of free $\mathcal{C}$-space, using the distance query from the workspace (Fig. 1). This enables fast $\mathcal{C}$-space exploration and path planning, which facilitates the algorithm operation in DEs. Roughly speaking, the generalized bur resembles the set of range sensors able to perceive the free $\mathcal{C}$-space locally and thus aid the algorithm in computing quality collision-free paths. This feature turns out to be particularly helpful in replanning stages. In the sequel, all components of DRGBT will be briefly presented referring to Fig. 2. In [50], the reader can find more technical details.
@@
DRGBT starts with planning of initial path from $\boldsymbol{q}_{start}$ to $\boldsymbol{q}_{goal}$ using any static planner. Specifically, we use RGBT-Connect [51] or RGBMT* [55]. If the path is found, it is labeled predefined path, which is then attempted to follow. Precisely, its $N_{h}$ nearest nodes are exploited to render a so-called local horizon $\boldsymbol{Q}_{h}$ (e.g., nodes $\boldsymbol{q}_{3},\dots,\boldsymbol{q}_{6}$ in Fig. 2 (a)). In case when the initial path is not found, $N_{h}$ random nodes are generated from the neighborhood of $\boldsymbol{q}_{start}$, which are then added to $\boldsymbol{Q}_{h}$. After the horizon is generated, horizon spines (i.e., generalized bur) are computed, starting from $\boldsymbol{q}_{curr}$ towards all nodes from $\boldsymbol{Q}_{h}$ to obtain a so-called reached horizon $\boldsymbol{Q}_{h}^{*}$ (as depicted in Fig. 2 (c)).
@@
The next step is the quality assessment of each reached node from $\boldsymbol{Q}_{h}^{*}$, where each one is assigned a node weight (e.g., the size of the blue circle centered at the corresponding node indicates the value of its weight in Fig. 2 (a), (b)). This procedure is a heuristic attempt to capture two important aspects: node proximity to the goal, and the minimal workspace distance to obstacles $d_{c}$, along with its rate. Furthermore, all estimated node weights are used to decide whether to replan the predefined path. If yes (e.g., Fig. 2 (b)), the replanning attempts to obtain a new path from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{goal}$. In case the new path is not found, the replanning is automatically triggered in the next iteration, while the robot structurally explores the neighborhood of $\boldsymbol{q}_{curr}$.
@@
$N_{h}=\min\left\{\left\lfloor N_{h_{0}}\left(1+\frac{d_{crit}}{d_{c}}\right)\right\rfloor,~{}n\cdot N_{h_{0}}\right\},\vspace{-0.15cm}$ (1)
@@
Finally, the node with the highest weight from $\boldsymbol{Q}_{h}^{*}$ is chosen as the next target node $\boldsymbol{q}_{next}$ at which the robot is headed (depicted as red “$\times$” in Fig. 2 (a), (b)). In case there are many nodes with the same weight, the closest one to $\boldsymbol{q}_{goal}$ is chosen.
II-C Anytime Feature of DRGBT
@@ II-D The Course of the Algorithm
The pseudocode of the algorithm is given in Alg. 1. Besides standard inputs required by DRGBT from [50], whose goal is to obtain collision-free geometric paths, the updated algorithm aims at computing time-parameterized trajectories and therefore has to account for constraints. To this end, we consider kinematic constraints $\mathcal{K}$ imposed on joint variables via maximal velocity $\boldsymbol{\omega}_{max}$, acceleration $\boldsymbol{\alpha}_{max}$, and (if available) jerk $\boldsymbol{j}_{max}$. The next input is real-time information about the environment, stored in $\mathcal{WO}$. The update of $\mathcal{WO}$ occurs within the function $\mathtt{perception}$ by providing the information/message about the actual pose/size of each obstacle. It is assumed that $\mathtt{perception}$ is executed by another processor in hard real-time, and, however important, is not a constitutive part of our planning algorithm. As for DRGBT’s outputs, the variable $goal\_reached$ denotes whether $\boldsymbol{q}_{goal}$ is reached, and $\boldsymbol{Q}_{trav}$ is a sequence containing visited robot’s configurations.
@@
Input: $\boldsymbol{q}_{start}$, $\boldsymbol{q}_{goal}$, $\mathcal{K}$, $\mathcal{WO}$, $safe\_on$
@@
Output: $goal\_reached$, $\boldsymbol{Q}_{trav}$
@@ 36             $goal\_reached\leftarrow\textnormal{{{false}}}$
41      if $replanning$ or $\mathtt{whetherToReplan}(\boldsymbol{w})$ then
@@
45            if $\boldsymbol{Q}_{pred,new}\,\neq\,\emptyset$ then

### v2

@@ Abstract
In this paper, we present the main features of Dynamic Rapidly-exploring Generalized Bur Tree (DRGBT) algorithm, a sampling-based planner for dynamic environments. We provide a detailed time analysis and appropriate scheduling to facilitate a real-time operation. To this end, an extensive analysis is conducted to identify the time-critical routines and their dependence on the number of obstacles. Furthermore, information about the distance to obstacles is used to compute a structure called dynamic expanded bubble of free configuration space, which is then utilized to establish sufficient conditions for a guaranteed safe motion of the robot while satisfying all kinematic constraints. An extensive comparative study is conducted to compare the proposed algorithm to competing state-of-the-art methods. Finally, an experimental study on a real robot is carried out covering a variety of scenarios including those with human presence. The results show the effectiveness and feasibility of real-time execution of the proposed motion planning algorithm within a typical sensor-based arrangement, using cheap hardware and sequential architecture, without the necessity for GPUs or heavy parallelization.
@@ I Introduction
TABLE I: Summary of the relevant features of the selected set of the state-of-the-art algorithms.
@@
replanning
freq. $\mathrm{[Hz]}$
@@ 2.60 GHz CPU
RRT-based algorithms for DEs have been concurrently developed with those based on PRM. One of the first approaches is execution-extended RRT (ERRT) algorithm [14], which improves replanning efficiency and the quality of generated paths by using a so-called waypoint cache memory in $\mathcal{C}$-space. Afterward, dynamic RRT (DRRT) algorithm [15], based on the D* family of deterministic algorithms [16], is proposed. It simply removes occupied edges in a tree, while collision-free ones are retained. Multipartite RRT (MP-RRT) algorithm [17] combines the advantages of existing RRT adaptations for dynamic MP from [18, 9, 15]. This is conveniently achieved by affecting a sampling distribution and reusing previously planned edges. Chance-constrained RRT [19] is an RTMP algorithm known for using so-called chance constraints to guarantee probabilistic feasibility in DE.
@@
Hardware acceleration has long been used in MP. The authors in [43] developed a dedicated FPGA-based circuitry to solve the planning problem in less than $\mathrm{1\,[ms]}$ for a 6-DoF manipulator. The collision detection time for PRM edges is demonstrated to be virtually independent of the graph size. Similarly, a massive GPU acceleration has been applied in [44] demonstrating a parallel optimization technique to solve the MP problem for manipulators. A significant speedup over state-of-the-art methods is achieved. GPU-accelerated library – CuRobo is released, which offers a parallel geometric planner and a collision-free IK solver. However, a price to pay in both [43] and [44] is having a dedicated, expensive hardware with a heavy GPU acceleration or specific FPGA-based circuitry.
@@
We stress that all of the aforementioned planners deal with at least one issue from the list below:
  • the algorithm is validated using low-DoF scenarios (e.g., mobile robots or robotic arms with fewer than six DoFs);
@@
To this end, we propose a novel approach that aims at overcoming all of the aforementioned issues. Relevant features of the state-of-the-art algorithms are summarized within Tab. I. The sign “n/a” stands for the feature that was either not available or not possible to find despite our best efforts.
@@
  • a comprehensive comparative study revealing DRGBT outperforms competing state-of-the-art methods;
@@
The remainder of the paper is structured as follows. Sec. II provides an overview of DRGBT algorithm from [50]. Sec. III proposes a scheduling framework for the real-time execution of the algorithm. Sec. IV proposes a concept of dynamic bubbles and burs, and proves that a guaranteed safe motion of the robot in DEs can be achieved. A simulation study within Sec. V reveals the most critical routines, and confirms that the chosen scheduling framework ensures the real-time execution of the algorithm. Moreover, a comparative study within Sec. VI juxtaposes DRGBT to state-of-the-art methods. Sec. VII provides experimental results demonstrating the real-time execution. Finally, Sec. VIII brings some discussion, concluding remarks, and future work directions.
Figure 1: Generalized bur with 30 “spines” (radial directions) and 5 “layers” (extensions along a single spine) (right) corresponding to a configuration of 2-DoF robot with four world obstacles (left) [51].
Figure 2: Graphical interpretation of some DRGBT components: obstacles (black); predefined/traversed path (gray/green); local (reached) horizon $\boldsymbol{Q}_{h}$ ($\color[rgb]{1,0,0}\boldsymbol{Q}_{h}^{*}$) with the corresponding horizon spines, i.e., generalized bur (solid red lines); next target node (red $\times$”); the value of node weight (the size of the blue circle centered at the corresponding node). (a) The obstacle blocks the predefined path to the goal; (b) The obstacle has moved in a way that clears the path to the goal, thus replanning has been performed and the predefined path has been changed; (c) The process of generating new local horizon and path replanning (as needed) continues.
@@ II DRGBT Algorithm
For better comprehension, we first introduce the notation and define the problem statement in Subsec. II-A. The background used within the proposed approach is described in Subsec. II-B. Then, we broadly describe DRGBT algorithm with its main features through Subsecs. II-C and II-D, while the details including pseudocode are given in Subsec. II-E.
@@ II-A Problem Statement and Assumptions
The robot’s $n$-dimensional configuration space is denoted as $\mathcal{C}$. The obstacle space $\mathcal{C}_{obs}\subseteq\mathcal{C}$ is the closed subset of $\mathcal{C}$ which implies the collision with obstacles or self-collision. The free space $\mathcal{C}_{free}=\mathcal{C}\backslash\mathcal{C}_{obs}$ is the open subset of $\mathcal{C}$ that the robot can reach. It is assumed that $\mathcal{C}_{obs}$ corresponds to a finite number $N_{obs}$ of (possibly overlapping) convex world obstacles $\mathcal{WO}_{j}$, $j\in\{1,2,\dots,N_{obs}\}$, existing in the robot’s workspace $\mathcal{W}$. The case of non-convex obstacles can be tackled whether by using some explicit decomposition techniques (e.g., [52]), or indirectly through algorithms for collision/distance checking that use an implicit representation of the environment via bounding volume hierarchies (BVHs), which typically rely on convex geometric primitives [53].
@@
A motion trajectory $\boldsymbol{\pi}(t)=\boldsymbol{\pi}[\boldsymbol{q}_{0},\boldsymbol{q}_{f}]$ represents a curve defined by a continuous mapping $\boldsymbol{\pi}:t\in[t_{0},t_{f}]\mapsto\mathcal{C}$ such that $\boldsymbol{\pi}(t_{0})\mapsto\boldsymbol{q}_{0}$ and $\boldsymbol{\pi}(t_{f})\mapsto\boldsymbol{q}_{f}$, where $t_{0}$ is the initial time and $t_{f}$ is a final time (in general not a priori given). A trajectory is considered valid if and only if both $\boldsymbol{\pi}(t)\cap\mathcal{C}_{obs}(t)=\varnothing$, $\forall t\in[t_{0},t_{f}]$, and it is feasible for the robot to follow $\boldsymbol{\pi}(t)$ within the limits of its kinematic and other constraints $\mathcal{K}$.
@@
II-B Background
The concept of bubbles of free $\mathcal{C}$-space was introduced in [54] as a method for defining collision-free regions around a given configuration based on a single distance information from the workspace. A bubble $\mathcal{B}(\boldsymbol{q},d_{c})$ at configuration $\boldsymbol{q}$ is computed using the minimal distance $d_{c}$ between the robot and the set of obstacles $\mathcal{WO}$ within the workspace $\mathcal{W}$.
Although bubbles of free $\mathcal{C}$-space have proven to be a valuable tool for path modification and motion planning, their underlying definition is rather conservative. To satisfy convexity constraints and ensure computational simplicity, the size of these bubbles is often unnecessarily restricted. In response to this limitation, a so-called complete bubble of free $\mathcal{C}$-space is proposed in [55]. Such alternative volume-based, less conservative structure still relies on the same input information as the original bubble. By intersecting $N$ rays originating from the configuration $\boldsymbol{q}$ with the complete bubble, so-called burs of free $\mathcal{C}$-space are acquired, as depicted by red lines in Fig. 1 (right).
@@ II-B Background
DRGBT algorithm, originally proposed in [50], uses the idea of generalized burs of free $\mathcal{C}$-space from [51]. This kind of bur is proven to capture relatively large portions of free $\mathcal{C}$-space, using the distance query from the workspace (Fig. 1). This enables fast $\mathcal{C}$-space exploration and path planning, which facilitates the algorithm operation in DEs. Roughly speaking, the generalized bur resembles the set of range sensors able to perceive the free $\mathcal{C}$-space locally and thus aid the algorithm in computing quality collision-free paths. This feature turns out to be particularly helpful in replanning stages.
II-C A Brief Overview of the Algorithm
@@
In the sequel, all components of DRGBT will be briefly presented referring to Fig. 2. The reader can find more technical details in [50]. DRGBT starts with planning of initial path from $\boldsymbol{q}_{start}$ to $\boldsymbol{q}_{goal}$ using any static planner. Specifically, we use RGBT-Connect [51] or RGBMT* [56]. If the path is found, it is labeled predefined path, consisting of nodes $\boldsymbol{q}_{1}=\boldsymbol{q}_{start},\,\dots,\boldsymbol{q}_{K}=\boldsymbol{q}_{goal}$, which is then attempted to follow. Precisely, its $N_{h}$ nearest nodes are exploited to render a so-called local horizon $\boldsymbol{Q}_{h}=\{\boldsymbol{q}_{k},\,\dots,\,\boldsymbol{q}_{k+N_{h}-1}\}$, $k\in\{2,\dots,K-N_{h}+1\}$ (e.g., nodes $\boldsymbol{q}_{3},\dots,\boldsymbol{q}_{6}$ in Fig. 2 (a)). If $K-N_{h}+1<k\leq K$, then $N_{h}-K+k$ random nodes, chosen from neighborhood of $\boldsymbol{q}_{curr}$, are added to the horizon in order to ensure the horizon size maintains $N_{h}$.
In case when the initial path is not found, $N_{h}$ random nodes are generated from the neighborhood of $\boldsymbol{q}_{start}$, which are then added to $\boldsymbol{Q}_{h}$. After the horizon is generated, horizon spines (i.e., generalized bur) are computed, starting from $\boldsymbol{q}_{curr}$ towards all nodes from $\boldsymbol{Q}_{h}$ to obtain a so-called reached horizon $\boldsymbol{Q}_{h}^{*}=\{\boldsymbol{q}_{k}^{*},\,\dots,\,\boldsymbol{q}_{k+N_{h}-1}^{*}\}$ (as Fig. 2 (c) shows).
@@
The next step is the quality assessment of each reached node from $\boldsymbol{Q}_{h}^{*}$, where each one is assigned a node weight (as depicted in Fig. 2). This procedure is a heuristic attempt to capture two important aspects: node proximity to the goal, and the minimal workspace distance to obstacles $d_{c}$, along with its rate. Furthermore, all estimated node weights are used to decide whether to replan the predefined path. If yes (e.g., Fig. 2 (b, c)), the replanning attempts to obtain a new path from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{goal}$. In case the new path is not found, the replanning is automatically triggered in the next iteration, while the robot structurally explores the neighborhood of $\boldsymbol{q}_{curr}$.
@@
$N_{h}=\min\left\{\left\lfloor N_{h_{0}}\left(1+\frac{d_{crit}}{d_{c}}\right)\right\rfloor,~n\cdot N_{h_{0}}\right\},\vskip-4.26773pt$ (1)
@@
Finally, the node with the highest weight from $\boldsymbol{Q}_{h}^{*}$ is chosen as the next target node $\boldsymbol{q}_{next}$ at which the robot is headed (as depicted by red $\times$” in Fig. 2). In case there are many nodes with the same weight, the closest one to $\boldsymbol{q}_{goal}$ is chosen.
II-D Anytime Feature of DRGBT
@@ II-E The Course of the Algorithm
The pseudocode of the algorithm is given in Alg. 1. Besides standard inputs required by DRGBT from [50], whose goal is to obtain collision-free geometric paths, the updated algorithm aims at computing time-parameterized trajectories and therefore has to account for constraints. To this end, we consider kinematic constraints $\mathcal{K}$ imposed on joint variables via maximal velocity $\boldsymbol{\omega}_{max}$, acceleration $\boldsymbol{\alpha}_{max}$, and (if available) jerk $\boldsymbol{j}_{max}$. The next input is real-time information about the environment, stored in $\mathcal{WO}$. The update of $\mathcal{WO}$ occurs within the function $\mathtt{perception}$ by providing the information/message about the actual pose/size of each obstacle. It is assumed that $\mathtt{perception}$ is executed by another processor in hard real-time, and, however important, is not a constitutive part of our planning algorithm. As for DRGBT’s outputs, the variable goal_reachedgoal\_reached denotes whether $\boldsymbol{q}_{goal}$ is reached, and $\boldsymbol{Q}_{trav}$ is a sequence containing visited robot’s configurations.
@@
Input: $\boldsymbol{q}_{start}$, $\boldsymbol{q}_{goal}$, $\mathcal{K}$, $\mathcal{WO}$, safe_onsafe\_on
@@
Output: goal_reachedgoal\_reached, $\boldsymbol{Q}_{trav}$
@@ 36     $goal\_reached\leftarrow\textnormal{{{false}}}$
41 if replanningreplanning or $\mathtt{whetherToReplan}(\boldsymbol{w})$ then
@@
45    if $\boldsymbol{Q}_{pred,new}\,\neq\,\varnothing$ then